# Structural fluctuations in active glasses

### v2

@@ VI Conclusion
In this work, we conducted a comprehensive analysis of the dynamics of active glasses using harmonic description. Previously, Henke et al. [45] have studied the velocity fields of ABP glasses using harmonic description to analyze confluent cells. Our work builds on this by considering three different active glass models, analyzing the displacement fields, and discussing the impact of the characteristic normal modes in glasses. These results will be useful for analyzing the displacement fields in a wide range of active glasses.
Towards the end of this paper, we would like to suggest a possible extension of our work. Firstly, we find it interesting to explore the structural fluctuations in active glasses with lower densities. The present work focused on densities above the jamming point, where each particle is always in contact with other particles. However, when the density is below the jamming point, the system remains in an active glass state while each particle is caged due to frequent collisions with other particles, as in the case of hard-sphere glasses. The dynamics of such active glasses are highly complex, and the harmonic description cannot be applied directly. Therefore, studying the structural fluctuations of such lower-density active glasses would be an important and fascinating topic. Secondly, we suggest studying the structural relaxations of active glasses. In thermal glasses, the structural relaxation has been discussed in terms of the excitation of quasi-localized vibrational modes. It is highly non-trivial to understand how the active forces excite these modes and trigger the rearrangement of particles. The present work provides a solid foundation to study such non-linear dynamics of active glasses.

### v3

@@ VI Conclusion
In this work, we conducted a comprehensive analysis of the dynamics of active glasses using harmonic description. Previously, Henke et al. [47] have studied the velocity fields of ABP glasses using harmonic description to analyze confluent cells. They obtained Eq. (43) for excitational energy, similar equation for Eq. (48) for mean square average of velocity. They also obtained Eq. (58) for correlation length of velocity fields although our Eq. (58) is for displacement fields. Our work builds on this by considering three different active glass models, analyzing the displacement fields, and discussing the impact of the characteristic normal modes in glasses. These results will be useful for analyzing the displacement fields in a wide range of active glasses.
Towards the end of this paper, we would like to suggest a possible extension of our work. Firstly, we find it interesting to explore the structural fluctuations in active glasses with lower densities. The present work focused on densities above the jamming point, where each particle is always in contact with other particles. However, when the density is below the jamming point, the system remains in an active glass state while each particle is caged due to frequent collisions with other particles, as in the case of hard-sphere glasses [59, 60, 61]. The dynamics of such active glasses are highly complex, and the harmonic description cannot be applied directly. Therefore, studying the structural fluctuations of such lower-density active glasses would be an important and fascinating topic. Secondly, we suggest studying the structural relaxations of active glasses. In thermal glasses, the structural relaxation has been discussed in terms of the excitation of quasi-localized vibrational modes [43]. It is highly non-trivial to understand how the active forces excite these modes and trigger the rearrangement of particles. The present work provides a solid foundation to study such non-linear dynamics of active glasses.